I need help figuring 5x + 8 + 3x - x + 5 = 6x - 3
Answer by vidhyak(98)

5x + 8 + 3x - x + 5 = 6x - 3
5x + 8 + 3x - x + 5 - 6x + 3 = 0
x + 16 = 0
x = -16